# Closed Orbits of Dynamically Convex Reeb Flows: Towards the HZ- and Multiplicity Conjectures

### v2

@@ 5.2.1. General observations
Our next goal is to ensure that $k_{ijl}\to\infty$ as $l\to\infty$ for all $i$ and $j$. To this end, we need to require that $\epsilon<1$ and $\sigma<\min a_{ij}$. (Recall that by the assumptions of the theorem $a_{ij}>0$ for all $i$ and $j$.) First, note that the sequence $\vec{k}_{l}$ is necessarily unbounded; for otherwise the entire sequence $K_{l}$ would be bounded due to (5.15). Furthermore, $\vec{k}_{l}\neq 0$ since otherwise we would have $K_{l}=0$ by again (5.15). (Here we use the fact that $10<\epsilon<$ 1.) In particular, for every $l$ at least one of the integers $k_{ijl}\neq 0$. Then by the second inequality in (5.14), for a fixed $l$, all $k_{ijl}$ are non-zero and have the same sign when $\sigma<\min a_{ij}$. Therefore, replacing $\vec{k}_{l}$ by $-\vec{k}_{l}$ when necessary, we can guarantee that $k_{ijl}>0$ for all $i$, $j$ and $l$. Clearly, after this sign change the sequence still has bounded gap. Finally, since the norm of $\vec{k}_{l}$ goes to infinity, we must have $k_{ijl}\to\infty$ for all $i$ and $j$ again by the second inequality in (5.14).

### v3

@@ 5.2.2. Preliminary observations
Our next goal is to ensure that $k_{ijl}\to\infty$ as $l\to\infty$ for all $i$ and $j$. To this end, we need to require that $\epsilon<1$ and $\sigma<\min a_{ij}$. (Recall that by the assumptions of the theorem $a_{ij}>0$ for all $i$ and $j$.) First, note that the sequence $\vec{k}_{l}$ is necessarily unbounded; for otherwise the entire sequence $K_{l}$ would be bounded due to (5.15). Furthermore, $\vec{k}_{l}\neq 0$ since otherwise we would have $K_{l}=0$ by again (5.15). (Here we use the fact that $0<\epsilon<1$.) In particular, for every $l$ at least one of the integers $k_{ijl}\neq 0$. Then by the second inequality in (5.14), for a fixed $l$, all $k_{ijl}$ are non-zero and have the same sign when $\sigma<\min a_{ij}$. Therefore, replacing $\vec{k}_{l}$ by $\vec{k}_{l}$ when necessary, we can guarantee that $k_{ijl}>0$ for all $i$, $j$ and $l$. Clearly, after this sign change the sequence still has bounded gap. Finally, since the norm of $\vec{k}_{l}$ goes to infinity, we must have $k_{ijl}\to\infty$ for all $i$ and $j$ again by the second inequality in (5.14).